OpenStudy (anonymous):

for f(x)=1/x-5 and g(x)=x^2+2 find:

OpenStudy (anonymous):

1. (f o g)(x)

OpenStudy (tkhunny):

You have written 1/x-5 = \(\dfrac{1}{x}-5\) Is this as you intended?

OpenStudy (anonymous):

The x-5 is together, so it would be \[\frac{ 1 }{ x-5 }\]

OpenStudy (tkhunny):

Parentheses will help: 1/(x-5) What is f(2)?

OpenStudy (mathmale):

Enclosing your x=5 in parentheses would remove the ambiguity. This is important in the long run, as it saves everyone time and effort when no clarification is needed.

OpenStudy (anonymous):

I just realized I used the wrong problem I was having trouble with, The one I needed help with was f(x) = 1/(x-5) and g(x) = x^2+2: find (g o f)(6)

OpenStudy (tkhunny):

(f o g)(6) = f(g(6)) It looks like you need g(6).

OpenStudy (anonymous):

This is how to problem is laid out, I'm just not too sure what to do. i keep thinking to plug my 6 into the equation for X, but I don't think that's the right way to go about it

OpenStudy (anonymous):

This is like, my first time seeing this stuff, so I'm a little lost is all

OpenStudy (tkhunny):

In that case, you need g(6). Go! What do you get for g(6)?

OpenStudy (anonymous):

38?

OpenStudy (tkhunny):

Perfect. What do you get for f(g(6)) = f(38)?

OpenStudy (anonymous):

Do I plug the 38 into 1/(x-5)?

OpenStudy (tkhunny):

That's what f(38) is. Go!